# COMMET: orders-of-magnitude speed-up in finite element method via batch-vectorized neural constitutive updates

### v1

@@ A.2.1 Invariants
The standard invariants, and the corresponding second and fourth order tensors as defined in 22 are given by
@@ Appendix C Data generation and NCM training
When training all NCMs used in this work, we follow the EUCLID paradigm for unsupervised discovery of material behaviour. More specifically, we use the NN-EUCLID framework of Thakolkaran et al. [21, 6]. In brief, this allows for the training of NCMs using full-field displacements and global reaction forces, both of which are physically obtainable from real experiments by using a combination of digital image correlation (DIC) and a load cell, i.e. stress measurements are not required. Given the known displacements and reaction forces $R^{\beta,t}$ on $\beta=1,...,n_{\beta}$ constrained boundaries at $t=1,...,n_{t}$ time steps, the parameters $\mathcal{Q}$ for a given NCM are obtained using
@@
Figure 15: Comparison of trained NCM behavior with ground truth. The NCMs are able to discover the ground truth behaviour accurately in all cases apart from biaxial tension.
The NCMs are able to discover the ground truth behaviour accurately in all cases apart from biaxial tension, the loading for which is outside of the training data.
@@ Appendix D Machine details
For completeness, we provide the details of the CPUs used in the computational experiments conducted in this work. The details of the CPU in the workstation and on the HPC nodes are provided in Tab.  1.
@@
Data and code will be made available after publication.

### v2

@@ A.2.1 Invariants
The standard invariants, and the corresponding second and fourth order tensors as defined in 20 are given by
@@ Appendix C Data generation and NCM training
When training all NCMs used in this work, we follow the EUCLID paradigm for unsupervised discovery of material behavior. More specifically, we use the NN-EUCLID framework of Thakolkaran et al. [21, 6]. In brief, this allows for the training of NCMs using full-field displacements and global reaction forces, both of which are physically obtainable from real experiments by using a combination of digital image correlation (DIC) and a load cell, i.e. stress measurements are not required. Given the known displacements and reaction forces $R^{\beta,t}$ on $\beta=1,...,n_{\beta}$ constrained boundaries at $t=1,...,n_{t}$ time steps, the parameters $\mathcal{Q}$ for a given NCM are obtained using
@@
All NCMs made use of a kinematic layer that maps the deformation gradient to polyconvex terms as follows:
$\mathcal{K}\left(\bm{F}\right)=\begin{bmatrix}\tilde{I_{1}}-3&\tilde{I_{2}}^{3/2}-3^{3/2}&(J-1)^{2}\end{bmatrix}\,$. (66)
The hyperparameters for the inner networks are provided in Tab.  1 and are comparable to those found in literature, e.g. [21, 6, 5].
Table 1: Hyperparameters used in NCMs
Parameter Value
MICNN [9, 21, 58, 59]:
    Number of hidden layers 3
    Number of neurons in each hidden layer 16
    Total number of learnable parameters 675
CANN [8, 5, 38, 60]:
    Exponents used in power layer $\{1,2\}$
    Terms used in function layer $\{f(x)=x,\quad f(x)=\text{exp}\left(x\right)\}$
    Total number of learnable parameters 33
ICKAN [6, 61]:
    Number of hidden layers 2
    Order of splines 4
    Number of grid points 30
    Total number of learnable parameters 508
@@
Figure 15: Comparison of trained NCM behavior with ground truth. The NCMs are able to discover the ground truth behavior accurately in all cases apart from biaxial tension.
The NCMs are able to discover the ground truth behavior accurately in all cases apart from biaxial tension, the loading for which is outside of the training data.
@@ Appendix D Machine details
For completeness, we provide the details of the CPUs used in the computational experiments conducted in this work. The details of the CPU in the workstation and on the HPC nodes are provided in Tab.  2.
@@
Appendix E The effect of model size
The hyperparameter choices and model sizes considered in this work are comparable to those commonly reported in the literature [21, 6, 5]. However, applications may require models with significantly more or fewer parameters than are typically employed. Consequently, it is also of interest to assess how the performance gains arising from vectorization, batching, and CGO depend on the size of the NCM.
To this end, we repeat the material point benchmarks presented in Section  4.1.2 for MICNNs with varying numbers of hidden layers and layer widths, thereby spanning a wide range of model sizes. In particular, MICNNs are constructed with 2, 3, and 4, hidden layers and layer widths of 2, 4, 8, 16, 32, and 64. In addition, to evaluate whether similar performance gains can be achieved for traditional constitutive laws, we repeat the material point benchmarks for a Gent–Thomas material model.
The resulting speed-ups are shown in Fig.  16 as a function of batch size for different NCM sizes. Overall, speed-ups due to vectorization, batching, and CGO are larger for smaller NCMs. We speculate that this is because the parameters for the NCM are held in cache during these computations. Hence, for bigger NCMs there is less cache memory available for the data on which the NCM operates. However, since the contents of the CPU cache cannot be directly inspected, this hypothesis cannot be conclusively verified. Unsurprisingly, the speed-up for the Gent-Thomas model is similar to that observed for NCMs with a small number of parameters, i.e. over three orders of magnitude. Hence, batching and vectorization also leads to significant performance gains for traditional constitutive models.
Figure 16: Effect of NCM size on speed-up due to batching, vectorization, and CGO: (left) combined speed-ups from CGO and vectorization relative to a non-vectorized, non-CGO baseline; (middle) speed-ups due to vectorization only; and (right) additional speed-ups due to CGO only at different batch sizes. While increasing model size reduces the attainable speed-up, improvements of more than two orders of magnitude are still achieved for models substantially larger than those typically reported in the literature.
Despite the observed decrease in speed-up with increasing model size, the performance gains remain substantial. In particular, for NCMs with approximately 13,000 parameters, speed-ups exceeding two orders of magnitude are still obtained. This is notable given that typical NCMs reported in the literature contain on the order of 1,000 parameters [21].
In addition to computational performance, memory requirements are an important practical consideration. The RAM usage as a function of batch size and NCM size is therefore shown in Fig.  17. For the largest NCM considered, containing approximately 13,000 parameters, the total RAM usage remains below 1 GB for batch sizes smaller than $10^{4}$. Notably, this range also encompasses the batch sizes that yield near-optimal speed-ups (Fig.  16). Such memory requirements are well within the capabilities of modern computing hardware. Furthermore, a comparison of the left and right columns of Fig.  17 clearly indicates that the use of CGO reduces overall RAM consumption.
Figure 17: Effect of NCM size on RAM usage: RAM consumption (left) with CGO and (right) without CGO, shown (a) as a function of the number of parameters and (b) as a function of batch size.
@@ Appendix E The effect of model size
The COMMET codebase is publicly available through https://commet-code.github.io/ and https://doi.org/10.5281/zenodo.17310682.